OpenStudy (anonymous):

Consider the arithmetic sequence 15, 13.2, 12, .... (a) Find the nth term of the sequence. >>>13.5+1.5n??<<< (b) Find the number of +ve terms. (c) What is the sum of all the +ve terms of the sequence? @ganeshie8 @hartnn @hba

hartnn (hartnn):

is that 15,13.5,12 .. ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how about b?

hartnn (hartnn):

n'th term needs to be positive, >0 13.5+1.5n >0 solve this for n

hartnn (hartnn):

or solve for 13.5 +1.5n = 0 to get which term =0 and then, one number less will be number of positive terms

OpenStudy (anonymous):

1.5n>-13.5 -ve number??

hartnn (hartnn):

oops, 13.5 + 1.5 n <= 0 to get which term will be non positive. then one less than n will give you number of positive terms

OpenStudy (anonymous):

wait for me, i have dinner first

hartnn (hartnn):

i am just trying to find which terms starts to become negative for that i equate n'th term formula with <=0 that will be first negative term

hartnn (hartnn):

okk..

hartnn (hartnn):

oh and your n'th term will have negative sign because d = -1.5 (not +1.5) 15+(n-1) (-1.5) = 15 +1.5 -1.5n = 16.5 -1.5 n

hartnn (hartnn):

now do 16.5-1.5n >0 to get number of positive terms

hartnn (hartnn):

you would know why d=-1.5, right ?

OpenStudy (anonymous):

yes, since it is decreasing n>9